# Enhancing the Efficiency of Time-Dependent Density Functional Theory Calculations of Dynamic Response Properties

### v1

@@ 3 Constraints-based noise attenuation
The convergence of TDDFT results with respect to $\eta$ in the imaginary time domain indicates that one can write the TDDFT data for the DSF as a sum of a physical part of the DSF, $S_{p}(\mathbf{q},\omega)$, and the quasi-periodic noise around it, $S_{n}(\mathbf{q},\omega)$, i.e.,
@@
The frequency of the quasi-periodic noise in the DSF data can be easily identified by examining the spectrum of the derivative of DSF, $\partial S(q,\omega)/\partial\omega$. This is illustrated in the top panels of Fig. 3. Panel (a) displays $\partial S(q,\omega)/\partial\omega$, while panel (b) shows the corresponding power spectrum. From Fig. 3(b), we observe that the noise spectrum contains prominent narrowband components, with two major bands highlighted using red marks along the frequency axis. Therefore, the task is to filter narrow-band noise from the reference DSF data while ensuring a high degree of agreement with the reference ITCF. In general, when identifying the frequencies of narrowband noise, complications that might be introduced by sharp DSF peaks associated with bound-state excitations can be circumvented by restricting the analysis of $\partial S(\mathbf{q},\omega)/\partial\omega$ to a relatively featureless region of the reference spectrum and use the smoothness of the derivative to impose a minimal stationary point condition for the filtering technique.
@@
Figure 3: (a) Derivative of the reference DSF, $\partial S(q,\omega)/\partial\omega$, with $\eta=0.1~{\rm eV}$ and at the same parameters as shown in Fig. 1. Panel (b) shows the power spectral density of $\partial S(q,\omega)/\partial\omega$ with the Fourier frequencies of the dominant narrowband noise oscillations marked on the x-axis. Panel (c) displays the grid search over multiple window-size-polynomial-degree-combinations (color map) and the optimum window size for each polynomial degree (black symbols). In panel (d) we compare the reference DSF with $\eta=0.1~{\rm eV}$ with the data obtained using the SG filer with different polynomial degrees and corresponding to them optimal window sizes as depicted in panel (a).
@@
Noise filtering is a common procedure in data processing in fields such as analytical chemistry, geosciences, and medicine (see, for example, Refs. 76, 77). Among the large variety of filters readily available for application, we use the Savitzky-Golay (SG) filter 78, which is commonly used in spectroscopy. The parameters of the SG filter are the polynomial order, which defines the degree of the polynomial fit, and the window length, which specifies the number of data points in the moving window. The choice of the optimal window length that attenuates narrow-band noises at a given polynomial degree can be performed automatically using various model selection metrics (scoring algorithms). We use the corrected Akaike Information Criterion (AICc) 79 to select filter parameters, which balances how well the filter fits the data with a penalty for unnecessary complexity. To address the specific concerns related to our application, we have introduced two additional penalties. The first is a frequency-domain penalty designed to prevent solutions from retaining contributions around specified target frequencies, which correspond to the dominant components of narrowband noise. The second is a second-derivative penalty that increases with the local curvature of the filtered signal. The application of this metric is illustrated in panel (c) of Fig. 3, which displays the grid search involving different combinations of window sizes and polynomial degrees, as well as the optimal window size for each polynomial degree. In panel (d), the DSFs obtained using the identified optimal SG filter parameters are compared to the reference DSF data. The data shown in panel (d) reveal that the noise-attenuated results for $S_{p}(\mathbf{q},\omega)$ closely match each other, demonstrating the effectiveness and reliability of the applied strategy.
@@
The requirement of minimizing the effect of filtering on the ITCF curvature, along with the condition outlined in Eq. (5), ensures that the frequency moments of the DSF are impacted as little as possible, since these moments can be expressed as the derivatives of the ITCF with respect to the imaginary-time argument $\tau$ 68, 58, 80.
@@
The close alignment of the obtained curves and their strict compliance with the specified constraints in the imaginary time domain indicate that the attenuation of narrow-band noise produces a reliable representation of $S_{p}(\mathbf{q},\omega)$. In summary, regardless of the noise-filtering strategy employed, the key criteria are (1) efficient attenuation of narrowband noise and (2) minimal impact on the ITCF.
Figure 5: Workflow for applying the $\eta$-convergence test in the imaginary time domain combined with the constraints-based noise attenuation technique. For the first step, the generation of the DSFs with various $\eta$ is realized at the post-processing phase of the TDDFT simulation and requires negligible computational time.
We stress that the use of the constraints-based filtering method extends beyond modeling the DSF, as the latter is directly related to the imaginary component of the dynamic density response function via the fluctuation-dissipation theorem 81:
@@
where $n$ is the electronic number density. Once $S_{p}(\mathbf{q},\omega)$ is determined, the fluctuation-dissipation theorem thus directly provides the imaginary part of the corresponding dynamic density response function. Furthermore, the Kramers-Kronig relations enable the computation of its real part 81 and corresponding complete knowledge of $\chi(\mathbf{q},\omega)$ gives one access to a host of dynamic dielectric properties, such as absorption coefficient and conductivity (see Sec. A.2). The workflow for applying the $\eta$-convergence test in the imaginary time domain, in conjunction with the constraints-based noise attenuation method, is summarized in Fig. 5.
Instead of starting from the DSF, one can also begin with any other dynamic dielectric property, e.g, conductivity. Then, by applying linear-response theory and the fluctuation-dissipation theorem, it is possible to derive the reference DSF data that are afflicted by the unphysical noise. Subsequently, the constraints-based filtering method can be used to determine an unbiased solution for $S_{p}(\mathbf{q},\omega)$. Using $S_{p}(\mathbf{q},\omega)$, one can then propagate backwards to compute smooth and physical curves for any relevant dynamic response properties.
@@ 4.1 Solid density hydrogen
Figure 6: (a) TDDFT results for the DSF of solid density hydrogen with $T=4.8~{\rm eV}$ and $\rho=0.08~{\rm g/cc}$ at $q=0.946~\textup{~\AA }^{-1}$. Panel (b) shows the corresponding ITCF results. In panel (c), we quantify the accuracy of the TDDFT results compared to exact PIMC data 59, 72 using the deviation measure (9). Shown is the difference between the DSF data computed using $\eta=0.1~{\rm eV}$ and the DSF results obtained using the constraints-based noise attenuation method and also using values of $\eta=0.3~{\rm eV}$ and $\eta=0.5~{\rm eV}$.
@@
As an independent check of the TDDFT data, we show the deviation of the TDDFT results for the ITCF from the exact path integral quantum Monte Carlo (PIMC) data  59 in Fig. 6(c), where the relative deviation is estimated using:
@@
where $\widetilde{F}_{\rm PIMC}(\mathbf{q},\tau)$ is the shifted ITCF data from PIMC. The TDDFT data with $\eta=0.1~{\rm eV}$ and the corresponding filtered results are in very good agreement with the exact PIMC results with $\Delta\widetilde{F}(\mathbf{q},\tau)<0.75~\%$. The growing deviation between the TDDFT and PIMC results as $\tau$ approaches $\beta/2$ is attributed to the adiabatic exchange correlation kernel approximation in our TDDFT simulations 73. As expected, the increase in $\eta$ to the values outside of the $\eta$-convergence range leads to an increase in the deviation of the TDDFT data from the exact PIMC results.
Figure 7: (a) Dynamic absorption coefficient (multiplied by the light speed) and (b) dynamic conductivity of solid density hydrogen with $T=4.8~{\rm eV}$ and $\rho=0.08~{\rm g/cc}$ at $q=0.946~\textup{~\AA }^{-1}$.
@@
In Fig. 6(d), we present the difference between the DSF data before and after applying the constraints-based noise elimination, i.e., $S_{n}(\mathbf{q},\omega)=S(\mathbf{q},\omega)-S_{p}(\mathbf{q},\omega)$. For comparison, in in Fig. 6(e) and in Fig. 6(f), we also show the difference between the DSF data computed using $\eta=0.1~{\rm eV}$ and the DSF results smoothed by increasing $\eta$ to $0.3~{\rm eV}$ and $0.5~{\rm eV}$. We observe that the constraints-based filtering produces $S_{n}(\mathbf{q},\omega)$ fluctuating symmetrically around zero. In contrast, the commonly used method of smoothing by increasing $\eta$ leads to a loss of physically relevant information, and, hence, biased results for the DSF.
@@
As discussed in Sec.3, the fluctuation-dissipation theorem and the Kramers-Kronig relations link the DSF with other dynamic response properties (see Sec. A.2). To illustrate the application of the $\eta$-convergence test in the imaginary time domain in conjunction with the constraints-based noise attenuation method, we present results for the absorption coefficient and dynamic conductivity of dense hydrogen in Fig. 7. We present the results based on $S_{p}(\mathbf{q},\omega)$, which were obtained through constraints-based filtering, along with the data generated from the reference DSF with $\eta=0.1~{\rm eV}$. Additionally, we compare to the results computed using $\eta=0.3~{\rm eV}$ and $\eta=0.5~{\rm eV}$. In Fig. 7, we see that the most notable discrepancies between the data based on $S_{p}(\mathbf{q},\omega)$ and the results obtained using $\eta=0.3~{\rm eV}$ and $\eta=0.5~{\rm eV}$ occur near their respective maximum values. Additionally, we note that the absorption coefficient is closely related to the opacity (see Sec. A.4), which is highly relevant for theories of stellar evolution and for practical applications such as IFE; see Ref. 82 and references therein.
@@
Figure 8: (a) TDDFT results for the DSF of Al with $T=0.025~{\rm eV}$ and for isochorically heated electrons with $T=25~{\rm eV}$. The electronic isosurfaces in fcc aluminum (boxes at the bottom of panel (a)) demonstrate the spread of electron density into the interstitial region as a result of isochoric heating 83. The bottom panels (b) and (c) display the $\eta$-convergence tests in the imaginary time domain using the minimum of the shifted ITCF at $\tau=\beta/2$ and the area under the shifted ITCF. The presented data are for the wavenumber $q=1.0~\textup{~\AA }^{-1}$ directed along the [100] direction relative to the crystal structure.
@@ 4.2 Aluminum with a face-centered cubic crystal structure
As a second example of applying the $\eta$-convergence test in the imaginary time domain, along with the constraints-based noise elimination method, we examine Al with a face-centered cubic (fcc) crystal structure. This material is frequently used as a testbed for new diagnostic techniques of electronic structure properties 84, 4 and for developing computational methods 85, 86, 26, 87.
@@
In Fig. 8(a), we present TDDFT results for the DSF of electrons in fcc aluminum with 11 valence electrons per atom at momentum transfer $q=1.0~\textup{~\AA }^{-1}$ along the [100] direction relative to the crystal structure. We analyze two cases: ambient conditions at a temperature of $T=0.025~\rm~{\rm eV}$ ($T\sim 290~{\rm K}$) and an isochorically heated state at $T=25~\rm~{\rm eV}$. The results of the analysis on the $\eta$-convergence in the imaginary time domain are shown in Figs. 8(b) and 8(c) for $T=0.025~\rm~{\rm eV}$ and for $T=25~\rm~{\rm eV}$, respectively, showing both the dependence on $\eta$ of the minimum of the shifted ITCF and the area under the shifted ITCF. We find that $\eta$-convergence is reached at $\eta\leq 0.5~{\rm eV}$. The DSF results computed with $\eta=0.5~{\rm eV}$ thus serve as the reference data for the application of the constraints-based SG filtering at $T=25~{\rm eV}$. In the isochorically heated scenario at a temperature of $T=25~\rm eV$, there is a significant computational bottleneck due to a large number of partially occupied orbitals and the requirement of considerable cutoff energy to describe thermal excitation from 2s and 2p states (L-shell) correctly. Considering the cubic crystal structure with a fcc crystal lattice parameter of $a=4.05\,\textup{~\AA }^{-1}$, we used 100 orbitals per atom and $2700~{\rm eV}$ cutoff energy to capture enough partially occupied KS orbitals and to describe the effect of thermal excitations; with the smallest occupation number being less than $10^{-5}$ and the $k$-point grid set to $10\times 10\times 10$ (see Sec. A.5 for the SG filter parameters). Here, we show the cold data to demonstrate the changes in the DSF due to thermal excitations. In the cold case, there are only six occupied KS orbitals per atom. With no computational issues, a noise-free DSF curve at $T=0.025~\rm~{\rm eV}$ was generated using a $40\times 40\times 40$ $k$-point grid. We emphasize that the one calculation with such a dense $k$-point grid at $T=25~\rm eV$ would have required about six million CPU hours (assuming the best case scenario with cubic scaling of the compute time with the density of $k$-point grid). For comparison, the presented approach combining the $\eta$-convergence test in the imaginary time domain and the constraints-based filtering allowed us to compute the noise-free DSF data using only $\sim 10^{5}$ CPU hours.
@@
In Fig. 8(a), one can see a significant broadening and shift of the plasmon to larger energies in the DSF at $T=25~\rm eV$ compared to the cold case due to strong thermal excitations. The thermal excitations also generate a prominent 2s-2p transition feature, as discussed in Ref. 88. In addition, we clearly see the L-shell features in both cold and hot cases. The reference DSF results at $T=25~\rm eV$ demonstrate significant fluctuations (noise) in both the plasmon and L-shell features. Increasing $\eta$ to $1.5~\rm eV$ reduces this noise, but it also leads to a strong damping of the bound-bound 2s-2p feature. In contrast, the constraints-based SG filtering effectively attenuates noise across the entire considered frequency range while preserving the 2s-2p feature.
@@
For $T=25~\rm eV$, we compare the data computed using $10\times 10\times 10$ and $18\times 18\times 18$ $k$-point grids (both with $\eta=0.5~{\rm eV}$). In addition, we show the results obtained by applying the constraints-based SG filtering to the data computed using $10\times 10\times 10$ $k$-point grid with $\eta=0.5~{\rm eV}$. At $T=25~\rm eV$, we observe that the constraints-based SG filtering allows us to obtain DSF data with the same quality as the simulations applying an $18\times 18\times 18$ $k$-point grid, where one can see a sharp and prominent 2s-2p feature and a broad Compton signal overlapping with the L-shell features. In this case, this constitutes more than a fivefold speedup, reducing the compute time from $\sim 8.4\times 10^{4}$ to $\sim 1.6\times 10^{4}$ CPU hours. In practice, to describe the experimental XRTS measurement for a specific scattering angle, it is necessary to perform calculations for several wavenumbers surrounding the central value in order to account for the finite size of the detector 4, 5. Additionally, the modeling of materials that are laser-heated or shock-compressed requires a series of simulations to consider variations in temperature and density within the target 89, and also potentially an even wider range of densities and temperatures for a statistical analysis of the most probable set of conditions 90. Consequently, the present combination of the $\eta$-convergence test in the imaginary time domain along with constraints-based noise attenuation has the potential for significant savings in computational time, and thereby to facilitate a more rigorous analysis of experimental measurements.

### v3

@@ 3 Constraints-based attenuation of narrow-band fluctuations
The convergence of TDDFT results with respect to $\eta$ in the imaginary time domain indicates that one can write the TDDFT data for the DSF as a sum of a physical part of the DSF, $S_{p}(\mathbf{q},\omega)$, and the quasi-periodic non-physical fluctuations around it, $S_{n}(\mathbf{q},\omega)$, i.e.,
@@
The frequency of the quasi-periodic fluctuations in the DSF data can be easily identified by examining the spectrum of the derivative of DSF, $\partial S(q,\omega)/\partial\omega$. This is illustrated in the top panels of Fig. 3. Panel (a) displays $\partial S(q,\omega)/\partial\omega$, while panel (b) shows the corresponding power spectrum. From Fig. 3(b), we observe that the spectrum of fluctuations contains prominent narrowband components, with two major bands highlighted using red marks along the frequency axis. Therefore, the task is to filter narrow-band fluctuations from the reference DSF data while ensuring a high degree of agreement with the reference ITCF. In general, when identifying the frequencies of narrowband fluctuations, complications that might be introduced by sharp DSF peaks associated with bound-state excitations can be circumvented by restricting the analysis of $\partial S(\mathbf{q},\omega)/\partial\omega$ to a relatively featureless region of the reference spectrum and use the smoothness of the derivative to impose a minimal stationary point condition for the filtering technique.
@@
Figure 3: (a) Derivative of the reference DSF, $\partial S(q,\omega)/\partial\omega$, with $\eta=0.1~{\rm eV}$ and at the same parameters as shown in Fig. 1. Panel (b) shows the power spectral density of $\partial S(q,\omega)/\partial\omega$ with the Fourier frequencies of the dominant narrowband fluctuations marked on the x-axis. Panel (c) displays the grid search over multiple window-size-polynomial-degree-combinations (color map) and the optimum window size for each polynomial degree (black symbols). In panel (d) we compare the reference DSF with $\eta=0.1~{\rm eV}$ with the data obtained using the SG filer with different polynomial degrees and corresponding to them optimal window sizes as depicted in panel (a).
@@
Filtering of non-physical fluctuations is a common procedure in data processing in fields such as analytical chemistry, geosciences, and medicine (see, for example, Refs. ACHARYA2016677, Schmid_ACS_2022). Among the large variety of filters readily available for application, we use the Savitzky-Golay (SG) filter SG_1964, which is commonly used in spectroscopy. The parameters of the SG filter are the polynomial order, which defines the degree of the polynomial fit, and the window length, which specifies the number of data points in the moving window. The choice of the optimal window length that attenuates narrow-band fluctuations at a given polynomial degree can be performed automatically using various model selection metrics (scoring algorithms). We use the corrected Akaike Information Criterion (AICc) AICC to select filter parameters, which balances how well the filter fits the data with a penalty for unnecessary complexity. To address the specific concerns related to our application, we have introduced two additional penalties. The first is a frequency-domain penalty designed to prevent solutions from retaining contributions around specified target frequencies, which correspond to the dominant components of narrow-band fluctuations. The second is a second-derivative penalty that increases with the local curvature of the filtered signal. The application of this metric is illustrated in panel (c) of Fig. 3, which displays the grid search involving different combinations of window sizes and polynomial degrees, as well as the optimal window size for each polynomial degree. In panel (d), the DSFs obtained using the identified optimal SG filter parameters are compared to the reference DSF data. The data shown in panel (d) reveal that the results for $S_{p}(\mathbf{q},\omega)$ with attenuated narrow-band fluctuations closely match each other, demonstrating the effectiveness and reliability of the applied strategy. The used postprocessing filter for the considered case is publicly available NoiseFilterTDDFT.
@@
The requirement of minimizing the effect of filtering on the ITCF curvature, along with the condition outlined in Eq. (5), ensures that the frequency moments of the DSF are impacted as little as possible, since these moments can be expressed as the derivatives of the ITCF with respect to the imaginary-time argument $\tau$ Dornheim_prb_2023, Dornheim_mre_2023, tolias2025exactseriesexpansionfrequency. It is important to note that a converged ITCF alone does not provide a unique determination of the DSF, since the inverse mapping (analytic continuation) from the ITCF in the imaginary-time domain to the DSF on the real-frequency axis is subject to practical numerical instabilities JARRELL1996133, PhysRevB.95.014102. Nevertheless, convergence of the ITCF is a necessary criterion for any reliable method for interpreting XRTS measurements at extreme conditions, as it automatically ensures convergence of a number of physical characteristics, including the frequency moments of the DSF Dornheim_prb_2023.
@@
The close alignment of the obtained curves and their strict compliance with the specified constraints in the imaginary time domain indicate that the attenuation of narrow-band fluctuations produces a reliable representation of $S_{p}(\mathbf{q},\omega)$. In summary, regardless of the filtering strategy employed, the key criteria are (1) efficient attenuation of narrowband fluctuations and (2) minimal impact on the ITCF.
Figure 5: Workflow for applying the $\eta$-convergence test in the imaginary time domain combined with the constraints-based attenuation of narrow-band fluctuations. For the first step, the generation of the DSFs with various $\eta$ is realized at the post-processing phase of the TDDFT simulation and requires negligible computational time.
We stress that the use of the constraints-based filtering method extends beyond modeling the DSF, as the latter is directly related to the imaginary component of the dynamic density response function via the fluctuation-dissipation theorem quantum_theory:
@@
where $n$ is the electronic number density. Once $S_{p}(\mathbf{q},\omega)$ is determined, the fluctuation-dissipation theorem thus directly provides the imaginary part of the corresponding dynamic density response function. Furthermore, the Kramers-Kronig relations enable the computation of its real part quantum_theory and corresponding complete knowledge of $\chi(\mathbf{q},\omega)$ gives one access to a host of dynamic dielectric properties, such as absorption coefficient and conductivity (see Sec. A.2). The workflow for applying the $\eta$-convergence test in the imaginary time domain, in conjunction with the constraints-based filtering of narrow-band fluctuations, is summarized in Fig. 5.
Instead of starting from the DSF, one can also begin with any other dynamic dielectric property, e.g, conductivity. Then, by applying linear-response theory and the fluctuation-dissipation theorem, it is possible to derive the reference DSF data that are afflicted by the unphysical narrow-band fluctuations. Subsequently, the constraints-based filtering method can be used to determine an unbiased solution for $S_{p}(\mathbf{q},\omega)$. Using $S_{p}(\mathbf{q},\omega)$, one can then propagate backwards to compute smooth and physical curves for any relevant dynamic response properties.
@@ 4.1 Solid density hydrogen
Figure 6: (a) TDDFT results for the DSF of solid density hydrogen with $T=4.8~{\rm eV}$ and $\rho=0.08~{\rm g/cc}$ at $q=0.946~\textup{~\AA }^{-1}$. Panel (b) shows the corresponding ITCF results. In panel (c), we quantify the accuracy of the TDDFT results compared to exact PIMC data Dornheim_MRE_2024, Bellenbaum_prr_2025 using the deviation measure (9). Shown is the difference between the DSF data computed using $\eta=0.1~{\rm eV}$ and the DSF results obtained using the constraints-based attenuation of narrow-band fluctuations and also using values of $\eta=0.3~{\rm eV}$ and $\eta=0.5~{\rm eV}$.
@@
As an independent check of the TDDFT data, we show the deviation of the TDDFT results for the ITCF from the exact path integral quantum Monte Carlo (PIMC) data  Dornheim_MRE_2024 in Fig. 6(c), where the relative deviation is estimated using:
@@
where $\widetilde{F}_{\rm PIMC}(\mathbf{q},\tau)$ is the shifted ITCF data from PIMC. The TDDFT data with $\eta=0.1~{\rm eV}$ and the corresponding filtered results are in very good agreement with the exact PIMC results with $\Delta\widetilde{F}(\mathbf{q},\tau)<0.75~\%$. The growing deviation between the TDDFT and PIMC results as $\tau$ approaches $\beta/2$ is attributed to the adiabatic exchange correlation kernel approximation in our TDDFT simulations moldabekov2025abinitiodensityfunctional. As expected, the increase in $\eta$ to the values outside of the $\eta$-convergence range leads to an increase in the deviation of the TDDFT data from the exact PIMC results.
Figure 7: Comparison of the DSF computed using $10\times 10\times 10$ and $20\times 20\times 20$ k-point grids for hydrogen at $T=4.8~\mathrm{eV}$ and $\rho=0.08~\mathrm{g/cc}$, at $q=0.946~\textup{\AA }^{-1}$. The results are shown for $\eta=0.2~\mathrm{eV}$ in panel (a) and for $\eta=0.1~\mathrm{eV}$ in panel (b).
@@
In Fig. 6(d), we present the difference between the DSF data before and after applying the constraints-based elimination of narrow-band fluctuations, i.e., $S_{n}(\mathbf{q},\omega)=S(\mathbf{q},\omega)-S_{p}(\mathbf{q},\omega)$. For comparison, in in Fig. 6(e) and in Fig. 6(f), we also show the difference between the DSF data computed using $\eta=0.1~{\rm eV}$ and the DSF results smoothed by increasing $\eta$ to $0.3~{\rm eV}$ and $0.5~{\rm eV}$. We observe that the constraints-based filtering produces $S_{n}(\mathbf{q},\omega)$ fluctuating symmetrically around zero. In contrast, the commonly used method of smoothing by increasing $\eta$ leads to a loss of physically relevant information, and, hence, biased results for the DSF.
The discussed quasi-periodic fluctuations are a direct consequence of the finite k-point grid. We demonstrate this by comparing data from calculations using $20\times 20\times 20$ and $10\times 10\times 10$ k-point grids for $\eta=0.1~{\rm eV}$ and $\eta=0.2~{\rm eV}$. In Fig. 7 (a), one can see that, at $\eta=0.2~{\rm eV}$, the DSF computed using the $20\times 20\times 20$ k-point grid (black dashed line) is substantially smoother than the results obtained with the $10\times 10\times 10$ grid, indicating a significant reduction of finite-k discretization effects. When the constraints-based filtering of narrow-band fluctuations is applied to the data generated with the $10\times 10\times 10$ grid, the resulting DSF (red solid line) closely matches the result computed using a $20\times 20\times 20$ k-point grid.
From panel (a) of Fig. 7, we observe that the reduction of $\eta$ from $\eta=0.2~\mathrm{eV}$ to $\eta=0.1~\mathrm{eV}$ leads to the appearance of quasi-periodic fluctuations even for the $20\times 20\times 20$ k-point grid, although with significantly smaller magnitude than in the $10\times 10\times 10$ case. This means that decreasing $\eta$ inevitably amplifies finite-k discretization effects for any finite k-point grid. Importantly, after applying the constraints-based procedure, the DSFs obtained from the two k-point grids (red solid line for $10\times 10\times 10$, orange dashed line for $20\times 20\times 20$) are smooth and in very close agreement with each other. This agreement indicates that, at $\eta=0.1~\mathrm{eV}$, the results obtained using the proposed approach are effectively converged with respect to k-point density.
Figure 8: (a) Imaginary part of the dynamic dielectric function and (b) dynamic conductivity of solid density hydrogen with $T=4.8~{\rm eV}$ and $\rho=0.08~{\rm g/cc}$ at $q=0.946~\textup{~\AA }^{-1}$.
@@
As discussed in Sec.3, the fluctuation-dissipation theorem and the Kramers-Kronig relations link the DSF with other dynamic response properties (see Sec. A.2). To illustrate the application of the $\eta$-convergence test in the imaginary time domain in conjunction with the constraints-based attenuation of narrow-band fluctuations, we present results for the Imaginary part of the dynamic dielectric function (which, e.g., describes electromagnetic wave absorption alexandrov1984principles) and dynamic conductivity of dense hydrogen in Fig. 8. We present the results based on $S_{p}(\mathbf{q},\omega)$, which were obtained through constraints-based filtering, along with the data generated from the reference DSF with $\eta=0.1~{\rm eV}$. Additionally, we compare to the results computed using $\eta=0.3~{\rm eV}$ and $\eta=0.5~{\rm eV}$. In Fig. 8, we see that the most notable discrepancies between the data based on $S_{p}(\mathbf{q},\omega)$ and the results obtained using $\eta=0.3~{\rm eV}$ and $\eta=0.5~{\rm eV}$ occur near their respective maximum values.
@@
Figure 9: (a) TDDFT results for the DSF of Al with $T=0.025~{\rm eV}$ and for isochorically heated electrons with $T=25~{\rm eV}$. The electronic isosurfaces in fcc aluminum (boxes at the bottom of panel (a)) demonstrate the spread of electron density into the interstitial region as a result of isochoric heating Moldabekov_jctc_2024. The bottom panels (b) and (c) display the $\eta$-convergence tests in the imaginary time domain using the minimum of the shifted ITCF at $\tau=\beta/2$ and the area under the shifted ITCF. The presented data are for the wavenumber $q=1.0~\textup{~\AA }^{-1}$ directed along the [100] direction relative to the crystal structure.
@@ 4.2 Aluminum with a face-centered cubic crystal structure
As a second example of applying the $\eta$-convergence test in the imaginary time domain, along with the constraints-based filtering of narrow-band fluctuations, we examine Al with a face-centered cubic (fcc) crystal structure. This material is frequently used as a testbed for new diagnostic techniques of electronic structure properties Preston_APL_2019, Gawne_PRB_2024 and for developing computational methods Panholzer_PRL_2018, Kononov2023, Witte_PRL, gawne2025orientationaleffectslowpair.
@@
In Fig. 9(a), we present TDDFT results for the DSF of electrons in fcc aluminum with 11 valence electrons per atom at momentum transfer $q=1.0~\textup{~\AA }^{-1}$ along the [100] direction relative to the crystal structure. We analyze two cases: ambient conditions at a temperature of $T=0.025~\rm~{\rm eV}$ ($T\sim 290~{\rm K}$) and an isochorically heated state at $T=25~\rm~{\rm eV}$. The results of the analysis on the $\eta$-convergence in the imaginary time domain are shown in Figs. 9(b) and 9(c) for $T=0.025~\rm~{\rm eV}$ and for $T=25~\rm~{\rm eV}$, respectively, showing both the dependence on $\eta$ of the minimum of the shifted ITCF and the area under the shifted ITCF. We find that $\eta$-convergence is reached at $\eta\leq 0.5~{\rm eV}$. The DSF results computed with $\eta=0.5~{\rm eV}$ thus serve as the reference data for the application of the constraints-based SG filtering at $T=25~{\rm eV}$. In the isochorically heated scenario at a temperature of $T=25~\rm eV$, there is a significant computational bottleneck due to a large number of partially occupied orbitals and the requirement of considerable cutoff energy to describe thermal excitation from 2s and 2p states (L-shell) correctly. Considering the cubic crystal structure with a fcc crystal lattice parameter of $a=4.05\,\textup{~\AA }^{-1}$, we used 100 orbitals per atom and $200~{\rm Ry}$ cutoff energy to capture enough partially occupied KS orbitals and to describe the effect of thermal excitations; with the smallest occupation number being less than $10^{-5}$ and the $k$-point grid set to $10\times 10\times 10$ (see Sec. A.4 for the SG filter parameters). Here, we show the cold data to demonstrate the changes in the DSF due to thermal excitations. In the cold case, there are only six occupied KS orbitals per atom. With no computational issues, a DSF curve free of non-physical fluctuations at $T=0.025~\rm~{\rm eV}$ was generated using a $40\times 40\times 40$ $k$-point grid.
@@
In Fig. 9(a), one can see a significant broadening and shift of the plasmon to larger energies in the DSF at $T=25~\rm eV$ compared to the cold case due to strong thermal excitations. The thermal excitations also generate a prominent 2s-2p transition feature, as discussed in Ref. baczewski2021predictionsboundboundtransitionsignatures. In addition, we clearly see the L-shell features in both cold and hot cases. The reference DSF results at $T=25~\rm eV$ demonstrate significant fluctuations in both the plasmon and L-shell features. Increasing $\eta$ to $1.5~\rm eV$ reduces these fluctuations, but it also leads to a strong damping of the bound-bound 2s-2p feature. In contrast, the constraints-based SG filtering effectively attenuates narrow-band fluctuations across the entire considered frequency range while preserving the 2s-2p feature.
@@
For $T=25~\rm eV$, we compare the data computed using $10\times 10\times 10$ and $32\times 32\times 32$ $k$-point grids (both with $\eta=0.5~{\rm eV}$). In addition, we show the results obtained by applying the constraints-based SG filtering to the data computed using $10\times 10\times 10$ $k$-point grid and $18\times 18\times 18$ $k$-point grid with $\eta=0.5~{\rm eV}$. At $T=25~\rm eV$, we observe that the constraints-based SG filtering using $10\times 10\times 10$ $k$-point grid allows us to obtain DSF data with the quality comparable to the simulations applying an $32\times 32\times 32$ $k$-point grid, where one can see a sharp and prominent 2s-2p feature and a broad Compton signal overlapping with the L-shell features. In this case, this constitutes more than a fifty-fold speedup, reducing the compute time from about $\sim 8.8\times 10^{5}$ to $\sim 1.6\times 10^{4}$ CPU hours ***This corresponds to worse than the ideal cubic scaling with respect to the k-point grid density, due to additional overhead associated with parallelization over 192 nodes with 96 CPUs per node.. In practice, to describe the experimental XRTS measurement for a specific scattering angle, it is necessary to perform calculations for several wavenumbers surrounding the central value in order to account for the finite size of the detector Gawne_PRB_2024, Gawne_2025. Additionally, the modeling of materials that are laser-heated or shock-compressed requires a series of simulations to consider variations in temperature and density within the target doi:10.1063/1.4893146, and also potentially an even wider range of densities and temperatures for a statistical analysis of the most probable set of conditions Kasim_POP_2019. Consequently, the present combination of the $\eta$-convergence test in the imaginary time domain along with constraints-based attenuation of narrow-band fluctuations has the potential for significant savings in computational time, and thereby to facilitate a more rigorous analysis of experimental measurements.